OpenStudy (anonymous):

How would I use implicit differentiation to derive y^4-2y=4x^3+x?

ganeshie8 (ganeshie8):

use chain rule : \(\large (y^4)' = 4y^{4-1} y' = 4y^3 y'\)

ganeshie8 (ganeshie8):

for your given equation, differentiate both sides with respect to x

ganeshie8 (ganeshie8):

\[\large y^4-2y=4x^3+x\] differentiate both sides with respect to x : \[\large 4y^3y'-2y'=4(3x^2)+1\]

ganeshie8 (ganeshie8):

isolate \(\large y'\) and you're done !

OpenStudy (anonymous):

thank you so much!